Solve for x:
(2x-5)/3-3/2 = (x+4)/5

Answer by checkley75(3666)

(2X-5)/3-1.5=(X+4)/5
(2X-5-1.5*3)/3=(X+4)/5
(2X-5-4.5)/3=(X+4)/5
(2X-9.5)/3=(X+4)/5 NOW CROSS MULTIPLY
3(X+4)=5(2X-9.5)
3X+12=10X-47.5
3X-10X=-47.5-12
-7X=-59.5
X=-59.5/-7
X=8.5 ANSWER
PROOF
(2*8.5-5)/3-1.5=(8.5+4)/5
(17-5)/3-1.5=12.5/5
12/3-1.5=2.5
4-1.5=2.5
2.5=2.5